OpenStudy (anonymous):

xy+y^2=1 ; find d^2y/dx^2

OpenStudy (anonymous):

use implicit differentiation

OpenStudy (anonymous):

how? can you teach me..

OpenStudy (anonymous):

@myko still there?

OpenStudy (anonymous):

find first derivative :\(\huge y'=y+xy'+2yy'\) or \(\huge y'=\frac{y}{1-x-2y}\) now differentiate implicitly again: \(\huge y''=\frac{y'(1-x-2y)-y(1-2y')}{(1-x-2y)^2}\) now substitute \(y'=\frac{y}{1-x-2y}\) in the last expretion

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

oh, I noticed, there is a sign mistake in expretion that i wrote for y''. The second term in nominator should be: +y(1+2y')

OpenStudy (anonymous):

you can check this link for more info on implicit differentiation: http://faculty.eicc.edu/bwood/math150supnotes/supplemental11.htm

OpenStudy (anonymous):

it's just what you need for this type of problem

OpenStudy (anonymous):

thanks :)